OpenStudy (anonymous):

〖lim〗(x→1 )⁡〖(x2-1)/(x-1)〗

OpenStudy (freckles):

It is looking a bit weird to me.

OpenStudy (freckles):

I will try to guess what I'm seeing... \[\lim_{x \rightarrow 1}\frac{x^2-1}{x-1}\]?

OpenStudy (freckles):

If that is the problem, here is a hint to the problem: Factor the top and then cancel a common factor.

OpenStudy (anonymous):

ok ((

OpenStudy (freckles):

what do you get when you factor x^2-1?

OpenStudy (anonymous):

i don't no

OpenStudy (freckles):

a^2-b^2=(a-b)(a+b)

OpenStudy (anonymous):

\[\left( \times +1 \right)\]

OpenStudy (freckles):

x+1 is what you get when the (x-1)'s cancel

OpenStudy (freckles):

So since x+1 is continuous at x=1 you can easily find the limit now but just plugging in 1

OpenStudy (anonymous):

(x+1)(x-1)/(x-1)= (x+1)

OpenStudy (freckles):

right

OpenStudy (freckles):

But you to take care of the limit part now

OpenStudy (anonymous):

yaay thank u (:

OpenStudy (freckles):

np